Trevor's Grout Company produces and sells grout to clients so they can install tile in their new homes. Trevor has supplied the following data about his grout company:

Kilograms (of grout) produced and sold 340,000
Sales revenue $ 1,910,000
Variable manufacturing expense $ 956,000
Fixed manufacturing expense $ 258,000
Variable selling and administrative expense $ 363,000
Fixed selling and administrative expense $ 224,000
Net operating income $ 109,000

Trevor's Grout Company's contribution margin ratio is closest to:

Garrison 16e Rechecks 2018-06-19

Multiple Choice

  • 81.0%

  • 30.9%

  • 74.8%

  • 49.9%
Answer #1

Variable costs = 956,000+363,000 = 1,319,000

Contribution margin ratio

= (Sales revenue - Variable costs)/Sales revenue

= (1,910,000-1,319,000)/1,910,000

= 30.9%